OpenStudy (anonymous):

Solving a quadratic - like equation... I am totally confused! Here is the problem: (x^2 + 4x)^2 - (x^2 + 4x) - 20=0. It says let A = X^2 + 4x, then A^2 - A - 20 =0. Then they factored it (A-5)(A+4) = 0 I understand how they factor it, but I do not understand how they got from step one to step two. Someone help me? Where did the A^2 - A - 20 = 0 come from?

hero (hero):

(x^2 + 4x)^2 - (x^2 + 4x) - 20=0 ( A )^2 - ( A ) - 20=0

hero (hero):

A^2 - A - 20 = 0

OpenStudy (anonymous):

Well, when you take (x^2 +4x)=A You do see that the question is (x^2 + 4x)^2 - (x^2 + 4x) - 20 =0 right? So just replace (x^2 +4x) with A, we get A^2 - A +20 =0 Then factorize it, as (A+4)(A-5) = 0 and you're all set. :)

hero (hero):

Anywhere you see x^2 + 4x, replace it with A

OpenStudy (anonymous):

oooooh duh!! Ok, that totally helps now. I don't know why I wasn't seeing that before.

hero (hero):

:D

OpenStudy (anonymous):

thank you, that cleared up so much confusion and frustration!

hero (hero):

yw :D